Please help solve by completing the square...
x^2 + 1/2x = 1

Answer by Theo(13342)

your equation is:

x^2 + (1/2)*x = 1

the coefficient of the x^2 term is 1.

the coefficient of the x term is (1/2)

take (1/2) of the coefficient of the x term to get (1/4)

square that to get (1/16)

your completing the square factor becomes:

(x + (1/4))^2 = 1 + (1/16)

take the square root of each side of this equation to get:

x + (1/4) = +/- sqrt (17/16)

subtract (1/4) from both sides of this equation to get:

x = -(1/4) +/- sqrt(17/16)

that should be your solution.

if it is, then you can plug those values of x into your original equation and the original equation should be true.

your values of x come out to be:


x = .780776406
x = -1.280776406

your original equation is:

x^2 + (1/2)*x = 1

plugging the first and second values of x respectively into that equation, you get:

1 = 1
1 = 1

this confirms that the solutions are good.